# The computational power of discrete chemical reaction networks with bounded executions

### v3

@@ 6.2 Impossibility of stably deciding majority and parity
In this section, we prove Theorem 6.9, which is a special case of our main negative result, Section 6.3. We give a self-contained proof of Theorem 6.9 because it is simpler and serves as an intuitive warmup to some of the key ideas used in proving Section 6.3, without the complexities of dealing with arbitrary semilinear sets.
Theorem 6.9 shows a limitation on the computational power of entirely execution bounded, all-voting CRNs, but it requires an additional constraint on the CRN for the result to hold (and we later give counterexamples showing that this extra hypothesis is provably necessary), described in the following definition.
@@
No non-collapsing, all-voting, entirely execution bounded CRD can stably decide the majority predicate $[X_{1}\geq X_{2}?]$ or the parity predicate $[X\equiv 1\mod 2?]$.
Proof 6.10.
@@
Since $\mathcal{D}$ is entirely execution bounded, by Theorem 6.5, $\mathcal{D}$ has a linear potential function $\Phi(\mathbf{x})=\mathbf{v}\cdot\mathbf{x}$, where $\mathbf{v}\geq\mathbf{0}$. Adding a single $X_{2}$ to $\mathbf{a}_{i}$ increases $\Phi$ by the constant $\mathbf{v}(X_{2})$. Since $\|\mathbf{a}_{i}\|$ grows without bound, the number of reactions to get from $\mathbf{a}_{i}+\{X_{2}\}$ to $\mathbf{b}_{i}$ increases without bound as $i\to\infty$, and since each reaction strictly decreases $\Phi$ by at least 1, the total change in $\Phi$ that results from adding $X_{2}$ and then going from $\mathbf{a}_{i}+\{X_{2}\}$ to $\mathbf{b}_{i}$ is unbounded in $i$, so unboundedly negative for sufficiently large $i$ (negative once $i$ is large enough that $\delta\|\mathbf{a}_{i}\|\geq\mathbf{v}(X_{2})+2$). Similarly, adding a single $X_{1}$ to $\mathbf{b}_{i}$ and going from $\mathbf{b}_{i}+\{X_{1}\}$ to $\mathbf{a}_{i+1}$, the resulting total change in $\Phi$ is unbounded and (for large enough $i$) negative.
@@
Theorem 6.9 is false without the non-collapsing hypothesis. The following collapsing, leaderless (but all-voting and entirely execution bounded) CRD stably decides majority: Species $X_{1},x_{1}$ vote yes, while $X_{2},x_{2}$ vote no:
@@
It has bounded executions from every configuration: $\min(\#X_{1},\#X_{2})$ of the first reaction can occur, and the other reactions decrease molecular count, so are limited by the total configuration size. However, it is collapsing since, for any $n$, there exists an input of size $n$ that reaches a stable configuration of size 1. Theorem 6.9 is similarly false without the all-voting hypothesis; for each of the reactions with one product above, add another non-voting product $W$. This converts the CRD to be non-collapsing but not all-voting. Of course, the execution bounded hypothesis is also necessary: the original population protocols paper [1] showed that all-voting, non-collapsing, leaderless population protocols can stably decide all semilinear predicates.
@@
[…]
full It has bounded executions from any configuration: exactly #$\#X_{1}+\#X_{0}-1$ reactions can occur since each reduces #$\#X_{1}+\#X_{0}$ by 1. Similar to above, by adding the non-voting product $W$ to each reaction above, this CRD becomes non-collapsing but not all-voting, showing that the all-voting hypothesis is also necessary for stably deciding parity.
@@ 6.3 Impossibility of stably deciding not eventually constant predicates
We now present our main negative result, Section 6.3, which generalizes Theorem 6.9 to show that such CRNs can stably decide only very limited (eventually constant) predicates.

### v4

@@ 6.2 Impossibility of stably deciding majority and parity
In this section, we prove Theorem˜6.7, which is a special case of our main negative result, LABEL:thm:non-collapsing_allvoting_entirely_execution_bounded_decide_almost_constant. We give a self-contained proof of Theorem˜6.7 because it is simpler and serves as an intuitive warmup to some of the key ideas used in proving LABEL:thm:non-collapsing_allvoting_entirely_execution_bounded_decide_almost_constant, without the complexities of dealing with arbitrary semilinear sets.
Theorem˜6.7 shows a limitation on the computational power of entirely execution bounded, all-voting CRNs, but it requires an additional constraint on the CRN for the result to hold (and we later give counterexamples showing that this extra hypothesis is provably necessary), described in the following definition.
@@
No non-collapsing, all-voting, entirely execution bounded CRD can stably decide the majority predicate $X_{1}\geq X_{2}?]$ or the parity predicate $X\equiv 1\mod 2?]$.
Proof 6.8.
@@
Since $\mathcal{D}$ is entirely execution bounded, by LABEL:thm:potential-function-iff-entirely-execution-bounded, $\mathcal{D}$ has a linear potential function $\Phi(\mathbf{x})=\mathbf{v}\cdot\mathbf{x}$, where $\mathbf{v}\geq\mathbf{0}$. Adding a single $X_{2}$ to $\mathbf{a}_{i}$ increases $\Phi$ by the constant $\mathbf{v}(X_{2})$. Since $\|\mathbf{a}_{i}\|$ grows without bound, the number of reactions to get from $\mathbf{a}_{i}+\{X_{2}\}$ to $\mathbf{b}_{i}$ increases without bound as $i\to\infty$, and since each reaction strictly decreases $\Phi$ by at least 1, the total change in $\Phi$ that results from adding $X_{2}$ and then going from $\mathbf{a}_{i}+\{X_{2}\}$ to $\mathbf{b}_{i}$ is unbounded in $i$, so unboundedly negative for sufficiently large $i$ (negative once $i$ is large enough that $\delta\|\mathbf{a}_{i}\|\geq\mathbf{v}(X_{2})+2$). Similarly, adding a single $X_{1}$ to $\mathbf{b}_{i}$ and going from $\mathbf{b}_{i}+\{X_{1}\}$ to $\mathbf{a}_{i+1}$, the resulting total change in $\Phi$ is unbounded and (for large enough $i$) negative.
@@
Theorem˜6.7 is false without the non-collapsing hypothesis. The following collapsing, leaderless (but all-voting and entirely execution bounded) CRD stably decides majority: Species $X_{1},x_{1}$ vote yes, while $X_{2},x_{2}$ vote no:
@@
It has bounded executions from every configuration: $\min(\#X_{1},\#X_{2})$ of the first reaction can occur, and the other reactions decrease molecular count, so are limited by the total configuration size. However, it is collapsing since, for any $n$, there exists an input of size $n$ that reaches a stable configuration of size 1. Theorem˜6.7 is similarly false without the all-voting hypothesis; for each of the reactions with one product above, add another non-voting product $W$. This converts the CRD to be non-collapsing but not all-voting. Of course, the execution bounded hypothesis is also necessary: the original population protocols paper [angluin2004computation] showed that all-voting, non-collapsing, leaderless population protocols can stably decide all semilinear predicates.
@@
It has bounded executions from any configuration: exactly $X_{1}+\#X_{0}-1$ reactions can occur since each reduces $X_{1}+\#X_{0}$ by 1. Similar to above, by adding the non-voting product $W$ to each reaction above, this CRD becomes non-collapsing but not all-voting, showing that the all-voting hypothesis is also necessary for stably deciding parity.
@@ 6.3 Impossibility of stably deciding not eventually constant predicates
We now present our main negative result, LABEL:thm:non-collapsing_allvoting_entirely_execution_bounded_decide_almost_constant, which generalizes Theorem˜6.7 to show that such CRNs can stably decide only very limited (eventually constant) predicates.